OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 6|6 – 4x| = 8x + 4

OpenStudy (anonymous):

6|6 – 4x| = 8x + 4 x = 1 x = 5/2 or x = 2 x = 5/2 x = 5/2or x = 1